### v1

@@ 5 Specification Tests
$H_{A}^{MC}:\mathrm{\tau}_{1,0}^{(l)}(t)\neq 0$
for some $t\in\{(l+1),\ldots,T\}$ and $l\in\{2,\ldots,T-1\}$. Note that the above tests can be carried out via moment equality tests. We will outline the test for $H_{0}^{MC}$ against $H_{A}^{MC}$ in what follows, but note that the same results hold for $H_{0}^{PT}$ against $H_{A}^{PT}$. In particular, we may use a statistic of the form:
$T_{1,0}^{s}=\sum_{(t,l)\in\mathcal{C}_{S}}\left(\sqrt{G}\widehat{\mathrm{\tau}}_{1,0}^{(l)}(t)\right)^{2},$ (16)
or
$T_{1,0}^{m}=\max_{(t,l)\in\mathcal{C}_{S}}\left|\sqrt{G}\widehat{\mathrm{\tau}}_{1,0}^{(l)}(t)\right|,$ (17)
where $\mathcal{C}_{S}$ is an index set $\mathcal{C}_{S}=\{(l,t):\;t\in\{(l+1),\ldots,T\};\;l\in\{2,\ldots,T-1\}\}$ with cardinality $\mathcal{C}_{S}$. For the second test, an analogous statistic can be used.
Remark 2: The tests above are designed to cover all pre-treatment periods provided there is only a finite number of them. However, in practice, as researchers may not want to use evidence from far away pre-treatment periods, we can restrict the number of pre-treatment periods used in the corresponding moment equality tests. At the same time, observe that the moment equality testing framework is more flexible than standard pre-trend tests that only test the significance of a (some) DiD parameter(s) relative to a single, fixed pre-treatment reference period. Specifically, the test statistics in (16) and (17) allow to include also comparisons of other combinations of pre-treatment periods. In practice, this may yield more power to detect specific violations of misclassification and PT in pre-treatment periods than what could be detected with only a fixed reference period.
Remark 3: If the researcher believes that also other pre-treatment periods are affected by misspecification, one may set a threshold, say $\kappa$, beyond which the researcher has certainty that no misclassification or anticipation is present so that the PT only test can be carried out.
@@
Since the limiting distribution of our tests under the null hypothesis is non-standard, we used the nonparametric (block) bootstrap to derive critical values. To do so, let
@@
[…]
$T_{1,0}^{s,b}=\sum_{(t,l)\in\mathcal{C}_{S}}\left(\sqrt{G}(\widehat{\mathrm{\tau}}_{1,0}^{(l),b}(t)-\widehat{\mathrm{\tau}}_{1,0}^{(l)}(t))\right)^{2},$
@@
$T_{1,0}^{m,b}=\max_{(t,l)\in\mathcal{C}_{S}}\left|\sqrt{G}(\widehat{\mathrm{\tau}}_{1,0}^{(l),b}(t)-\widehat{\mathrm{\tau}}_{1,0}^{(l)}(t))\right|.$
Here, $\widehat{\mathrm{\tau}}_{1,0}^{(l),b}(t)$ denotes the bootstrap analog of the test statistic:
$\widehat{\mathrm{\tau}}_{1,0}^{(l),b}(t)=\frac{1}{N_{1,0,t}^{b}}\sum_{g:D^{b}_{g,t}=1,D^{b}_{g,t-1}=0}N_{g,t}\left(Y_{g,t-1}^{b}-Y_{g,t-l}^{b}\right)-\frac{1}{N^{b}_{0,0,t}}\sum_{g:D^{b}_{g,t}=0,D^{b}_{g,t-1}=0}N_{g,t}\left(Y_{g,t-1}^{b}-Y_{g,t-l}^{b}\right).$
@@
Finally, note that by Assumptions 19 and 20, we have with probability approaching 1 that there exists at least one $(t,l)$ for which $\lim_{G\rightarrow\infty}\mathrm{E}\left[P_{1,0}(t,l)\right]>0$ and so $\min\{N_{1,0,t},{N}_{0,0,t}\}>0$ for sufficiently large $G$. On the other hand, observe that for any element $(t,l)\in\mathcal{C}_{S}$ for which $\lim_{G\rightarrow\infty}\mathrm{E}\left[P_{1,0}(t,l)\right]=0$, the corresponding element does not contribute to the statistic $T_{1,0}^{s}$ for sufficiently large $G$, and does also not contribute to the asymptotic variance of $\sqrt{G}\left(\widehat{\boldsymbol{\tau}}_{1,0}-\mathbf{0}\right)$. We therefore proceed by analyzing the case where all elements of $\sqrt{G}\left(\widehat{\boldsymbol{\tau}}_{1,0}-\mathbf{0}\right)$ contribute to the asymptotic distribution, i.e. the case where $\lim_{G\rightarrow\infty}\mathrm{E}\left[\frac{N_{1,0,t}}{G}\right]>0$ and $\lim_{G\rightarrow\infty}\mathrm{E}\left[\frac{N_{0,0,t}}{G}\right]>0$ hold for all $(t,l)\in\mathcal{C}_{S}$ in Assumption 20. We have the following statement:
@@
Assume that Assumptions 1-2, 4-7, and 14, 19, 20 hold for all $(t,l)\in\mathcal{C}_{S}$, and that $\boldsymbol{\Sigma}_{\mathbf{U}}$ from (22) is positive definite, and $\mathbf{J}_{f}\left(\mathbf{U}^{\infty}\right)$ defined in (27) has full column rank. Then:
@@
As remarked previously, we note that identical results for the test $H_{0}^{PT}$ against $H_{A}^{PT}$ as we also demonstrate in our Monte Carlo simulations.
Finally, even though both tests can be carried out idependently, we may also consider a decision rule for the joint execution of both tests that provides guidance for practicioners. In particular, there are three cases we would like to distinguish as they bear different implications for the empirical analysis (see above): a violation of PT in pre-treatment periods, possibly with misclassification ($H_{TV}$), PT in pre-treatment periods holding without misclassification ($H_{PTN}$), and PT in pre-treatment periods holding with misclassification ($H_{PTM}$). We have that:
@@ 6 Simulations
[…]
where $\Delta_{t}$ is the treatment effect, while $e_{g,t}$ denotes the idiosyncratic error term. We simulate both time constant and time-varying treatment effects. Specifically, the constant treatment effects are drawn from a normal distribution with a mean of 4 and a variance of 1, i.e. $\Delta\sim\mathcal{N}(4,1)$. The time varying treatment effects on the other hand are generated by
@@
Table 1 and Table 2 display the simulation results for the estimators presented in Section 4 in comparison to the estimator of De Chaisemartin and D’Haultefoeuille (2020) that does not account for misspecification in the treatment indicator. Table 1 contains the mean bias with respect to $\Delta$ or $\bar{\Delta}\equiv 1/14\sum_{t=2}^{15}\Delta_{t}$, respectively. [^12] Note that in population the benchmark ATE coincides with the ATE of true switchers since the misclassification or switching probabilities are constant over time in this simulation setting. Turning to the results, the mean biases in the third column of Table 1 are not close to zero reflecting that $\widehat{\mathrm{DID}}$ cannot match the ATE in the presence of misclassification. This is due to some actually treated groups (individuals) entering the comparison group that is supposed to be non treated. If treatment effects change over time, the inclusion of already treated groups in the reported pre-treatment periods is also an issue. As the fourth column shows, the estimator for observed switchers $\widehat{\mathrm{DID}}^{S}$ captures the ATE under constant treatment effects, but not under time-varying effects. Under constant treatment effects there are no wrong comparisons and the exact timing of the treatment does not matter since the effect size is the same in all periods. The mean bias of the estimator for true switchers $\widehat{\mathrm{DID}}^{S^{*}}$ is close to zero in all cases and decreases as the total number of groups increases.
Table 2 depicts the Root Mean Squared Error (RMSE). For the $\widehat{\mathrm{DID}}$ estimator, the RMSE is larger under constant treatment effects than under time varying treatment effects (see column 3), which reflects the same pattern seen for the mean bias. The RMSE of the estimator for true switchers $\widehat{\mathrm{DID}}^{S^{*}}$ is of similar magnitude or larger than those of the estimator for observed switchers $\widehat{\mathrm{DID}}^{S}$, even though the former do not exhibit a bias. This is to be expected since $\widehat{\mathrm{DID}}^{S^{*}}$ requires the estimation of unobserved switching probabilities. Not surprisingly, the estimator for true switchers $\widehat{\mathrm{DID}}^{S^{*}}$ produces a larger RMSE under time-varying treatment effects than under constant treatment effects given that the mean bias is close to zero in both cases.
@@
$\bf{\widehat{\mathrm{DID}}}$ $\bf{\widehat{\mathrm{DID}}^{S}}$ $\bf{\widehat{\mathrm{DID}}^{S^{*}}}$
G = 100 Constant TE -2.3332 0.0317 -0.0298
Time-vary. TE -2.0845 0.2536 -0.0432
G = 300 Constant TE -2.3287 -0.0029 -0.0020
Time-vary. TE -2.0786 0.2139 -0.0190
G = 600 Constant TE -2.3307 -0.0136 -0.0010
Time-vary. TE -2.0818 0.1981 -0.0098
Table 1: Results of 2,000 Monte Carlo replications for 100, 300 and 600 groups are presented. Columns 3-5 contain the simulation means of the difference between the estimates and the mean treatment effects. Column 3 refers to the De Chaisemartin and D’Haultefoeuille (2020) estimator $\widehat{\mathrm{DID}}$, column 4 to the DID estimator of observed switchers $\widehat{\mathrm{DID}}^{S}$ and column 5 to the DID estimator of true switchers $\widehat{\mathrm{DID}}^{S^{*}}$.
@@
$\bf{\widehat{\mathrm{DID}}}$ $\bf{\widehat{\mathrm{DID}}^{S}}$ $\bf{\widehat{\mathrm{DID}}^{S^{*}}}$
G = 100 Constant TE 2.4241 0.2050 0.2312
Time-vary. TE 2.1731 0.3835 0.8699
G = 300 Constant TE 2.4080 0.1119 0.1102
Time-vary. TE 2.1513 0.2713 0.2842
G = 600 Constant TE 2.4072 0.0820 0.0783
Time-vary. TE 2.1514 0.2312 0.1197
Table 2: Results of 2,000 Monte Carlo replications for 100, 300 and 600 groups are presented. Columns 3-5 contain the root mean squared errors (RMSE) between the estimates and the mean treatment effects. Column 3 refers to the De Chaisemartin and D’Haultefoeuille (2020) estimator $\widehat{\mathrm{DID}}$, column 4 to the DID estimator of observed switchers $\widehat{\mathrm{DID}}^{S}$ and column 5 to the DID estimator of true switchers $\widehat{\mathrm{DID}}^{S^{*}}$.

### v2

@@ 5 Specification Tests
$H_{A}^{MC}:\mathrm{\tau}_{1,0}(t,l)\neq 0$
for some $t\in\{(l+1),\ldots,T\}$ and $l\in\{2,\ldots,T-1\}$. As mentioned previously, the above tests may be implemented as moment equality tests. We outline the test statistic for $H_{0}^{MC}$ against $H_{A}^{MC}$ in what follows, but note that the statistic for $H_{0}^{PT}$ against $H_{A}^{PT}$ is analogous. Thus, depending on the preferences of the researcher, one may use a Carmér-von-Mises type statistic that is geared towards detecting many small violations of the null hypothesis:
$T_{1,0}^{s}=\sum_{(t,l)\in\mathcal{C}_{S}}\left(\sqrt{G}\widehat{\mathrm{\tau}}_{1,0}(t,l)\right)^{2},$ (16)
or a Kolmogorov-Smirnov type max statistic:
$T_{1,0}^{m}=\max_{(t,l)\in\mathcal{C}_{S}}\left|\sqrt{G}\widehat{\mathrm{\tau}}_{1,0}(t,l)\right|,$ (17)
where $\mathcal{C}_{S}$ is an index set $\mathcal{C}_{S}=\{(l,t):\;t\in\{(l+1),\ldots,T\};\;l\in\{2,\ldots,T-1\}\}$ with cardinality $\mathcal{C}_{S}$.
Remark 4: The tests above are designed to cover all pre-treatment periods provided there is only a finite number of them. However, in practice, as researchers may not want to use evidence from far away pre-treatment periods, we can restrict the number of pre-treatment periods used in the corresponding moment equality tests.
Remark 5: If the researcher believes that also other pre-treatment periods are affected by misspecification, one may set a threshold, say $\kappa$, beyond which the researcher has certainty that no misclassification or anticipation is present, so that the PT only test can be carried out accordingly.
Remark 6: Note that the moment equality testing framework may be used more broadly as a flexible and potentially more powerful alternative to standard pre-trend tests in event-study type analyses. That is, while standard tests only evaluate the significance of a (or some) pre-trend DiD parameter(s) relative to a single, fixed pre-treatment reference period, the statistics in (16) and (17) may be modified to include also comparisons of other pre-treatment period combinations. For instance, if there are three pre-treatment periods (t3)(t-3), (t2)(t-2), and (t1)(t-1) available in the data, the moment based test may be used to compare not only outcomes from period (t1)(t-1) to periods (t2)(t-2) and (t3)(t-3), but also outcomes from (t2)(t-2) to (t3)(t-3). In practice, this may render the moment based test more powerful to detect specific violations of misclassification or PT in pre-treatment periods than what would be possible if only a fixed reference period was used (see also Section 7).
@@
Since the limiting distribution of our tests under the null hypothesis is non-standard, we use the nonparametric (block) bootstrap to derive critical values. To do so, let
@@
To this end, denote by $\mathbf{Y}_{g}^{b}$ the bootstrapped observations. In addition, we write $o^{b}_{\Pr}(1)$, in probability when for any $\delta>0$, $\operatorname*{\text{Pr}^{b}}\left(|T^{b}|>\delta\right)=o_{\Pr}(1)$ for any bootstrap statistic $T^{b}$. Also, we write $T^{b}\stackrel{{\scriptstyle d^{b}}}{{\rightarrow}}D$ in probability, if, conditional on the sample, $T^{b}$ weakly converges to $D$ with probability going to one. Finally, let $\mathrm{E}^{b}\left[\cdot\right]$, $\text{Var}^{b}\left[\cdot\right]$, and $\text{Cov}^{b}\left[\cdot\right]$ denote the expectation, variance, and covariance, respectively, under the bootstrap probability measure $\operatorname*{\text{Pr}^{b}}$, conditional on the original sample (defined on a given probability space $(\Omega,\mathcal{F},\Pr)$). [^10] The probability law governing the resample is given by the random variables $I_{g}$, $g=1,\ldots,G$, where each $I_{g}$ is an i.i.d. discrete uniform random variable on the interval $[1,G]$. We draw $B$ bootstrap samples $\{\mathbf{Y}_{g}^{b},\mathbf{D}_{g}^{b}\}_{g=1}^{G}$. The bootstrap statistics are given by:
$T_{1,0}^{s,b}=\sum_{(t,l)\in\mathcal{C}_{S}}\left(\sqrt{G}(\widehat{\mathrm{\tau}}_{1,0}^{b}(t,l)-\widehat{\mathrm{\tau}}_{1,0}(t,l))\right)^{2},$
@@
$T_{1,0}^{m,b}=\max_{(t,l)\in\mathcal{C}_{S}}\left|\sqrt{G}(\widehat{\mathrm{\tau}}_{1,0}^{b}(t,l)-\widehat{\mathrm{\tau}}_{1,0}(t,l))\right|.$
Here, $\widehat{\mathrm{\tau}}_{1,0}^{b}(t,l)$ denotes the bootstrap analog of the test statistic:
$\widehat{\mathrm{\tau}}_{1,0}^{b}(t,l)=\frac{1}{N_{1,0,t}^{b}}\sum_{g:D^{b}_{g,t}=1,D^{b}_{g,t-1}=0}N_{g,t}\left(Y_{g,t-1}^{b}-Y_{g,t-l}^{b}\right)-\frac{1}{N^{b}_{0,0,t}}\sum_{g:D^{b}_{g,t}=0,D^{b}_{g,t-1}=0}N_{g,t}\left(Y_{g,t-1}^{b}-Y_{g,t-l}^{b}\right).$
@@
Note that by Assumptions 19 and 20, we have with probability approaching 1 that there exists at least one $(t,l)$ for which $\lim_{G\rightarrow\infty}\mathrm{E}\left[P_{1,0}(t,l)\right]>0$ and so $\min\{N_{1,0,t},{N}_{0,0,t}\}>0$ for sufficiently large $G$. On the other hand, observe that for any element $(t,l)\in\mathcal{C}_{S}$ for which $\lim_{G\rightarrow\infty}\mathrm{E}\left[P_{1,0}(t,l)\right]=0$, the corresponding element does not contribute to the statistic $T_{1,0}^{s}$ for sufficiently large $G$, and does also not contribute to the asymptotic variance of $\sqrt{G}\left(\widehat{\boldsymbol{\tau}}_{1,0}-\mathbf{0}\right)$. We therefore proceed by analyzing the case where all elements of $\sqrt{G}\left(\widehat{\boldsymbol{\tau}}_{1,0}-\mathbf{0}\right)$ contribute to the asymptotic distribution, i.e. the case where $\lim_{G\rightarrow\infty}\mathrm{E}\left[\frac{N_{1,0,t}}{G}\right]>0$ and $\lim_{G\rightarrow\infty}\mathrm{E}\left[\frac{N_{0,0,t}}{G}\right]>0$ hold for all $(t,l)\in\mathcal{C}_{S}$ in Assumption 20. We have the following statement:
@@
Assume that Assumptions 1-2, 6-7, and 14, 19, 20 hold for all $(t,l)\in\mathcal{C}_{S}$, and that $\boldsymbol{\Sigma}_{\mathbf{U}}$ from (22) is positive definite, and $\mathbf{J}_{f}\left(\mathbf{U}^{\infty}\right)$ defined in (27) has full column rank. Then:
@@
Proposition 7 establishes the asymptotic size control of the moment based tests under $H_{0}^{MC}$, and their consistency against fixed alternatives. As remarked previously, identical results also hold for the test of $H_{0}^{PT}$ against $H_{A}^{PT}$. Moreover, note that the asymptotic analysis of these nonparametric moment tests relies on less and substantially weaker assumptions than the analysis of the bias and the estimators, respectively. Specifically, apart from the condition about the timing of misclassification (Assumption 7) to make the tests operational, we do not impose any restrictions on the form of misclassification (e.g., homogeneity in Assumption 12), strong exogeneity (Assumption 4), or PT (Assumption 5). The tests therefore allow to assess misspecification under a substantially weakened set of assumptions.
Finally, even though both tests may of course be carried out separately, we can also combine them to construct a decision rule that provides guidance for practicioners. In particular, there are three cases we would like to distinguish as they bear different implications for the empirical analysis (see above): a violation of PT in pre-treatment periods, possibly with misclassification ($H_{TV}$), PT in pre-treatment periods holding without misclassification ($H_{PTN}$), and PT in pre-treatment periods holding with misclassification ($H_{PTM}$). We have that:
@@
Even though the formal analysis of this decision rule is beyond the scope of the paper, the heuristic is as follows: if there is pre-treatment evidence suggesting a likely violation of PT, DID should generally not be used for point identification and estimation. On the other hand, if there is pre-treatment evidence supporting PT, we may actually proceed with DID and seek to design the estimator in such a way to avoid misspecification bias. Put differently, we aim to “maximize” our confidence in the PT assumption in a first testing step by looking at pre-treatment evidence in favor of PT, before turning to a misspecification test in a second step to obtain guidance on how to manipulate the estimator to avoid bias. In fact, building on pre-treatment evidence for PT is in line with recent work by Rambachan2023 who exploit pre-treatment PT violations as a way to infer about possible post-treatment violations. Moreover, the decision rule is also motivated by findings in CH2024 who show that conducting pre-treatment specification tests generally leads to, under the null of correct specification, conservative, but not too liberal post-treatment inference. Applied to our context, we therefore conjecture that post-treatment inference carried out conditional on passing the above pre-treatment specification tests may also maintain asymptotic size control under the null of correct specification.
@@ 6 Simulations
Since our asymptotic theory operates at the group level, we generate group level outcomes instead of modeling individual outcomes and taking group averages. We use 2,000 Monte Carlo replications and let the total number of groups be $G\in\{100,\ 300,\ 600\}$ across 15 time periods ($T=15$). The first treatment period for each group is randomly drawn between time periods $t=2$ and $t=15$. A fraction of 5% of the groups remain untreated throughout the panel. The DGP is chosen to be:
$Y_{g,t}(D_{g,t}^{*})=10-0.4\,t+0.1\,g+\Delta_{t}\,\mathbf{1}\{D_{g,t}^{*}\geq 1\}+e_{g,t},\;e_{g,t}\sim\mathcal{N}(0,1),$
where $\Delta_{t}$ is the treatment effect, while $e_{g,t}$ denotes the idiosyncratic error term. We simulate both time constant and time-varying treatment effects. Specifically, the constant treatment effect $\Delta$ is set to 4 and the time varying treatment effects are generated by
@@
Table 1 and Table 2 display the simulation results for the estimators presented in Section 4 in comparison to the estimator of CH2020 that does not account for misspecification in the treatment indicator. As $\widehat{\mathrm{DID}}^{S}$ and $\widehat{\mathrm{DID}}^{S^{*}}$ recover different target parameters, Table 1 contains the mean bias with respect to $\bar{\Delta}_{S}*\equiv 1/13\sum_{t=2}^{14}\Delta_{t}=4$ and $\bar{\Delta}_{S}\equiv 1/12\sum_{t=3}^{14}\Delta_{t}=4.2287$, respectively. [^11] Note that in population the benchmark ATE $\bar{\Delta}_{S}$ coincides with the ATE of true switchers since misclassification and switching probabilities are constant over time in the DGP. Under constant treatment effects, both parameters are equivalent, i.e. $\Delta=\bar{\Delta}_{S}*=\bar{\Delta}_{S}$. Turning to the results, the mean biases in the third and fourth columns of Table 1 are not close to zero reflecting that $\widehat{\mathrm{DID}}$ cannot match either of the ATEs in the presence of misclassification. This is due to some actually treated groups (individuals) entering the comparison group that is supposed to be non treated. If treatment effects change over time, the inclusion of already treated groups in the reported pre-treatment periods is also an issue. As the fifth column shows, the estimator for observed switchers $\widehat{\mathrm{DID}}^{S}$ captures the average treatment effect between time periods 3 to 14. The mean bias of the estimator for true switchers $\widehat{\mathrm{DID}}^{S^{*}}$ is close to zero in all cases and decreases as the total number of groups increases. Since this estimator captures anticipatory treatment effects, treatment effects in the second time period are also included.
Table 2 depicts the Root Mean Squared Error (RMSE). For the $\widehat{\mathrm{DID}}$ estimator, the RMSE is larger under constant treatment effects than under time varying treatment effects (see column 3 and 4), which reflects the same pattern seen for the mean bias. For $G=100$ and $G=300$ the RMSE of the estimator for true switchers $\widehat{\mathrm{DID}}^{S^{*}}$ is larger than the RMSE of the estimator for observed switchers $\widehat{\mathrm{DID}}^{S}$. The estimator for true switchers $\widehat{\mathrm{DID}}^{S^{*}}$ produces a larger RMSE under time-varying treatment effects than under constant treatment effects. This is to be expected since $\widehat{\mathrm{DID}}^{S^{*}}$ requires the estimation of unobserved switching probabilities.
@@
$\bf{\widehat{\mathrm{DID}}}$ $\bf{\widehat{\mathrm{DID}}}$ $\bf{\widehat{\mathrm{DID}}^{S}}$ $\bf{\widehat{\mathrm{DID}}^{S^{*}}}$
G = 100 Constant TE -2.3324 -2.3324 -0.0040 -0.0296
Time-vary. TE -2.0799 -2.3086 -0.0015 -0.0373
G = 300 Constant TE -2.3402 -2.3402 0.0010 -0.0038
Time-vary. TE -2.0895 -2.3181 0.0007 -0.0165
G = 600 Constant TE -2.3406 -2.3406 0.0011 -0.0007
Time-vary. TE -2.0895 -2.3182 0.0024 -0.0062
Table 1: Results of 2,000 Monte Carlo replications on the Mean Bias for 100, 300 and 600 groups are presented. Columns 3 and 4 refer to the CH2020 estimator $\widehat{\mathrm{DID}}$ in comparison to $\bar{\Delta}_{S}*$ and $\bar{\Delta}_{S}$, respectively. Column 5 refers to the DID estimator of observed switchers $\widehat{\mathrm{DID}}^{S}$ in comparison to $\bar{\Delta}_{S}$, and column 6 to the DID estimator of true switchers $\widehat{\mathrm{DID}}^{S^{*}}$ in comparison to $\bar{\Delta}_{S}*$. Note that $\bar{\Delta}_{S}$ and $\bar{\Delta}_{S^{*}}$ are identical under constant treatment effects.
@@
$\bf{\widehat{\mathrm{DID}}}$ $\bf{\widehat{\mathrm{DID}}}$ $\bf{\widehat{\mathrm{DID}}^{S}}$ $\bf{\widehat{\mathrm{DID}}^{S^{*}}}$
G = 100 Constant TE 2.3527 2.3527 0.1908 0.2904
Time-vary. TE 2.1060 2.3321 0.2570 0.7409
G = 300 Constant TE 2.3468 2.3468 0.1087 0.1067
Time-vary. TE 2.0979 2.3257 0.1503 0.1956
G = 600 Constant TE 2.3441 2.3441 0.0766 0.0743
Time-vary. TE 2.0938 2.3221 0.1060 0.1112
Table 2: Results of 2,000 Monte Carlo replications on the RMSE for 100, 300 and 600 groups are presented. Columns 3 and 4 refer to the CH2020 estimator $\widehat{\mathrm{DID}}$ in comparison to $\bar{\Delta}_{S}*$ and $\bar{\Delta}_{S}$, respectively. Column 5 refers to the DID estimator of observed switchers $\widehat{\mathrm{DID}}^{S}$ in comparison to $\bar{\Delta}_{S}$, and column 6 to the DID estimator of true switchers $\widehat{\mathrm{DID}}^{S^{*}}$ in comparison to $\bar{\Delta}_{S}*$. Note that $\bar{\Delta}_{S}$ and $\bar{\Delta}_{S^{*}}$ are identical under constant treatment effects.